ness00:

A youth group is setting up camp. Rain is predicted, so the campers decide to build a fly, or rain cover, over their tent. The fly will be 12 feet high and 16 feet wide. The scouts are building the frame for the fly with two poles slanted and joined together at the top of the tent. The minimum length of the slanted poles needed to support the fly is feet.

OLIVER69:

So a fly is usually a rectangular shape. So if we know that then it will be easy to find the length of the slanted poles. And it's fairly easy to do, all you would have to do is do 2 times the width (16). which would give you 32 for the length of the pole

Narad:

Apply Pythagoras Length of pole \[=\sqrt{\left( 8^{2}+12^{2} \right)}=\sqrt{208}\]